OpenStudy (anonymous):

the 10th term of an arithematic series is 34, and the sum of the first 20 terms is 710. determine the 25th term

OpenStudy (mertsj):

\[a _{n}=a _{1}+(n-1)d\]

OpenStudy (mertsj):

\[34=a _{1}+(10-1)d\]

OpenStudy (anonymous):

dont u use the formula sn = n/2 (2a (n-1)d) For the sum of the arethematic SERIES, NOT SEQUESNCE

OpenStudy (anonymous):

@nupursehrawat get here.. and see how it goes ;) ... (we can chat there too. ) http://www.twiddla.com/807035

OpenStudy (anonymous):

tell me when you're there ...

OpenStudy (anonymous):

@nupursehrawat ... are you there?

OpenStudy (anonymous):

@nupursehrawat check this out ...

OpenStudy (mertsj):

She told you that you couldn't use that formula.

OpenStudy (mertsj):

That's what stumped me. I thought we needed that formula.

OpenStudy (anonymous):

now we have... \[\Large a_1=34-9\cdot 3\] \[\Large a_1=7\] Now substitute \[\LARGE a_1=7 \quad \quad d=3\] in this equation... \[\LARGE a_{25}=a_1+(n-1)d\]

OpenStudy (anonymous):

look again , she was confused, and asked you if we should use formula of Sn instead of a_n but she forgot this : \[\huge ?\]

OpenStudy (mertsj):

I guess you and I are the only ones still interested in the problem.

OpenStudy (anonymous):

100% agree !

OpenStudy (anonymous):

no im here

OpenStudy (anonymous):

My mom Called me SOOOOOOOOORY

OpenStudy (anonymous):

Ok then check out link I posted before, Hope that helps , if it doesn't I don't know other way !

OpenStudy (anonymous):

It Does Help Thank You Verryyyy MUCH Appreciate It :)

OpenStudy (anonymous):

My preasure ... @Mertsj could do this one easily I'm extremely sure, but he thought that you're not allowed to use the formula I used LOL ...

OpenStudy (anonymous):

But Its Hard To Understand

OpenStudy (anonymous):

What You Did In The Picture

OpenStudy (anonymous):

do you know two main formulas of arithmetic sequences ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you mind if I ask you to write them for me ?

OpenStudy (anonymous):

Sn = N/2 (2a (n-1)D Sn = (a + Tn) Tn = term number a = first number

OpenStudy (anonymous):

other formula ?

OpenStudy (anonymous):

Its ok I Will ask my teacjer for help Bye

OpenStudy (anonymous):

...as you wish !